OpenStudy (muzzack):

@Monster1234 i need help and @caleb_gajewsky

OpenStudy (anonymous):

yess....

OpenStudy (muzzack):

i have issues

OpenStudy (muzzack):

;(

OpenStudy (anonymous):

what kind? young man

OpenStudy (anonymous):

what are those problems ?

OpenStudy (muzzack):

lol jk

OpenStudy (anonymous):

hahahaa

OpenStudy (muzzack):

u just got online prank'd

OpenStudy (muzzack):

yeahyuh

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

lol

OpenStudy (muzzack):

okay my real question is

OpenStudy (anonymous):

CALL THE POLICE!!!

OpenStudy (muzzack):

wht is the formula for the axis of symmetry

OpenStudy (acxbox22):

\(\Huge\color{red}{(:0|)}\)

OpenStudy (anonymous):

The graph of a quadratic function is a parabola. The axis of symmetry of a parabola is a vertical line that divides the parabola into two congruent halves. The axis of symmetry always passes through the vertex of the parabola. The x-coordinate of the vertex is the equation of the axis of symmetry of the parabola.

OpenStudy (muzzack):

is it \(\frac{-a}{2b}\)

OpenStudy (anonymous):

Y=ax^2+bx+c

OpenStudy (anonymous):

oh wait for the vertical... its x=-b/2a

OpenStudy (muzzack):

kkkkkkkkkkkkkkkk thx

OpenStudy (anonymous):

lol I just googled that up there :(

OpenStudy (anonymous):

wellcome mickey mouse